I need some help on this problem.
Which is the better buy: 5 lb. of sugar for $4.75 or 20 lb. of sugar for $19.92?
Thanks,
Lyn

Found 2 solutions by Nate, stanbon:
Answer by Nate(3500)

$4.75/5 = cost per pound of sugar = $0.95
$19.92/20 = cost per pound of sugar = $0.996
The smaller sugar bag is the best buy.

Answer by stanbon(48558)

Which is the better buy: 5 lb. of sugar for $4.75 or 20 lb. of sugar for $19.92?
4.75/5 = 0.95 cents per lb.
19.92/20 = 1.992/2= 0.996 cents per lb.
Looks like the 5 lb bag is the better buy.
Cheers,
Stan H.